OpenStudy (el_arrow):

convert the polar equation to rectangular form r=4sin(theta)

OpenStudy (el_arrow):

how do you do this?

OpenStudy (el_arrow):

please explain

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is?

OpenStudy (el_arrow):

no....

OpenStudy (el_arrow):

yeah i do

OpenStudy (jdoe0001):

hmmm ok... what section is this exercise covering anyway?

OpenStudy (el_arrow):

is taking the square

OpenStudy (el_arrow):

polar coordinates

OpenStudy (jdoe0001):

perfect square trinomial \(\large { \begin{array}{cccccllllll} {\color{brown}{ a}}^2& + &2{\color{brown}{ a}}{\color{blue}{ b}}&+&{\color{blue}{ b}}^2\\ \downarrow && &&\downarrow \\ {\color{brown}{ a}}&& &&{\color{blue}{ b}}\\ &\to &({\color{brown}{ a}} + {\color{blue}{ b}})^2&\leftarrow \end{array}\qquad % perfect square trinomial, negative middle term \begin{array}{cccccllllll} {\color{brown}{ a}}^2& - &2{\color{brown}{ a}}{\color{blue}{ b}}&+&{\color{blue}{ b}}^2\\ \downarrow && &&\downarrow \\ {\color{brown}{ a}}&& &&{\color{blue}{ b}}\\ &\to &({\color{brown}{ a}} - {\color{blue}{ b}})^2&\leftarrow \end{array} }\) looks familiar?

OpenStudy (el_arrow):

its like a^2+2ab+b^22

OpenStudy (jdoe0001):

well.. more or less.. yes one sec

OpenStudy (el_arrow):

okay

OpenStudy (jdoe0001):

\(\bf r=4sin(\theta)\qquad \begin{cases} x=rcos(\theta)\\ y={\color{brown}{ rsin(\theta) }} \\ tan(\theta)=\cfrac{y}{x}\\ x^2+y^2={\color{blue}{ r^2}} \end{cases}\qquad thus \\ \quad \\ r\cdot r=r\cdot 4sin(\theta)\implies {\color{blue}{ r^2 }} =4{\color{brown}{ rsin(\theta)}} \implies x^2+y^2=4y\) so... now it's just a matter of "completing the square" to make it look like a conic equation

OpenStudy (el_arrow):

okay so the 4y would go to the left of the equation right?

OpenStudy (jdoe0001):

yes \(\bf x^2+y^2=4y\implies x^2+y^2-4y=0 \\ \quad \\ x^2+(y^2-4y+{\color{red}{ \square ?}}^2)=0\) so.. any ideas on what we're missing there to "complete the square"?

OpenStudy (el_arrow):

(-4/2)^2

OpenStudy (jdoe0001):

hmm well.. the middle term is 2 * "other terms" so let's see the middle term here is "4y" leaving the negative out ...hmm

OpenStudy (jdoe0001):

well.... -4/2 will be -2 but yes the 3rd term in a "perfect square trinomial" is positive though so we'll use +2 :) one sec

OpenStudy (el_arrow):

okay

OpenStudy (jdoe0001):

keep in mind that if we ADD \(2^2\) we also have to SUBTRACT \(2^2\) since all we're really doing, is borrowing from our good fellow Mr Zero, "0" thus \(\bf x^2+y^2=4y\implies x^2+y^2-4y=0 \\ \quad \\ x^2+(y^2-4y+{\color{red}{ 2}}^2)-{\color{red}{ 2}}^2=0\implies x^2+(y-2)^2-4=0 \\ \quad \\ x^2+(y-2)^2=4\implies (x-0)^2+(y-2)^2=2^2\)

OpenStudy (jdoe0001):

so... as you can see is just a circle with a center at (0,2) with a radius of 2

OpenStudy (el_arrow):

oh okay

OpenStudy (el_arrow):

so the rectangular for is (x-0)^2+(y-2)^2=4

OpenStudy (jdoe0001):

yeap

OpenStudy (el_arrow):

thank you

OpenStudy (jdoe0001):

yw